OpenStudy (anonymous):

N(x)=100+20*ln(0.25x) explanation in comments please look

OpenStudy (anonymous):

he number of computers sold by Bcc depends on dollar mountx, that they spend on advertising. how many computers will they sell by spending 40000 dollars

OpenStudy (anonymous):

284?

OpenStudy (kropot72):

Yes. The correct answer is 284 computers.

OpenStudy (anonymous):

@kropot72 N(x)=100+20*ln(0.25x). the number of computers sold by bcc depends on dollar amount x, that they spend on advertising. how much do they spend to get 200 computers? how would i do this one

OpenStudy (kropot72):

200 = 100 + 20 * ln(0.25x) 200 - 100 = 20 * ln(0.25x) \[\frac{200-100}{20}=\ln (0.25x)\] 5 = ln(0.25x) Can you solve it now?

OpenStudy (anonymous):

@kropot72 maybe im still a little confused

OpenStudy (kropot72):

5 = ln(0.25x) \[e ^{5}=0.25x\] \[x=\frac{e ^{5}}{0.25}\]

OpenStudy (anonymous):

@kropot72 593.34

OpenStudy (kropot72):

My calculator gives $593.65

OpenStudy (anonymous):

@kropot72 did you use 2.718 for e

OpenStudy (kropot72):

No. When I enter e^1 on my calculator it gives the value of e to 9 decimal places.

OpenStudy (anonymous):

ohh ok for my quwstion we are just required to use 2.718 @kropot72

OpenStudy (kropot72):

Understood. In that case your answer should be marked as correct :)

OpenStudy (anonymous):

298 is the correct answer for this problem.